# Towards chemical accuracy for chemi- and physisorption with an efficient density functional

### v1

@@ II.3 Determination of parameters for Opt(MS+rVV10)
Here, $\kappa$ and $b$ are tunable parameters in the MS2 functional, while $\beta$ is a parameter in the rVV10 functional. In the original formulation of the MS2 meta-GGA, $\kappa$ = 0.504 and $b$ = 4.0 were optimized to fit the atomization energies of six molecules (AE6 set), and the barrier heights of six reactions (BH6 set) [67]. Similarly in rVV10, $\beta$ = 6.3 was fitted to the interaction energies of 22 predominant dispersion-dominated complexes (S22 set) [68].
In this work, we re-optimized $\kappa$, $b$, and $\beta$ to improve the functional’s accuracy for both chemisorption and physisorption. Since the MS2 meta-GGA is known to partially capture intermediate-range vdW interactions [58], it was necessary to re-fit $\beta$ of rVV10 to avoid overestimating vdW forces, which could lead to double-counting. The $\beta$ parameter of rVV10 was therefore optimized for each ( $\kappa$, $b$ ) pair by fitting the binding energy curve of $Ar_{2}$, using highly accurate coupled-cluster CCSD(T) calculations as the reference [69]. The $\kappa$ and $b$ of MS2 was then re-optimized by fitting to the experimental adsorption energy of CO on Pt(111), a well-studied surface system with reliable experimental data [70].
We minimized the absolute error in the CO/Pt(111) adsorption energy with respect to $\kappa$ and b using a least-squares fit to a 3rd-order polynomial in the two-dimensional $\kappa$-$b$ space. This approach allowed us to efficiently explore the parameter space and identify the optimal values. The resulting heat map (Figure 2(a)) revealed a well-defined minimum at $\kappa$=0.25 and $b$=0.91, with $\beta$ =26.26. This set of parameters delivered the lowest error for the CO/Pt(111) adsorption energy.
When using AE6 atomization energies as the optimization objective for $\kappa$, and $b$, a very different minimum was found, located in the upper-right corner of the $\kappa$-$b$ parameter space (Figure 2(b)). A similar minimum emerged when both the AE6 and BH6 sets were used, in line with the original MS2 parameterization. These findings suggest that developing DFT exchange-correlation functionals for molecular adsorption on solid surfaces requires distinct optimization strategies compared to those used for general chemistry applications.
@@
We conducted additional testing of the newly optimized Opt(MS+rVV10) functional on a set of small benchmark systems. These tests included the atomization energies from the AE6 set [67], hydrogen transfer barrier heights from the BH6 set [67], and lattice constants from the LC20 set [71], comprising 20 solids. The performance of Opt(MS+rVV10) was compared against other van der Waals-corrected functionals, including RPBE+D3, BEEF-vdW, and SCAN+rVV10 (the latter being a widely regarded general-purpose functional). The results are summarized in Table 1. For AE6 atomization energies, SCAN+rVV10 and BEEF-vdW showed better accuracy than Opt(MS+rVV10), which exhibited slightly higher errors. For BH6 barrier heights, both BEEF-vdW and Opt(MS+rVV10) performed well, with mean absolute errors (MAEs) of 5.4 kcal/mol and 5.6 kcal/mol, respectively. In the case of LC20 lattice constants, Opt(MS+rVV10) and SCAN+rVV10 were the most accurate, with MAEs of 0.020 Å and 0.021 Å, respectively. These results demonstrate that Opt(MS+rVV10) performs competitively in terms of accuracy, particularly for surface adsorption, while still maintaining reasonable accuracy in other general-purpose tasks.
III Conclusion
@@
We present the Opt(MS+rVV10), a novel exchange-correlation vdW-corrected functional, representing a significant advancement in the modeling of molecular adsorption on solid surfaces. This functional achieves a balanced performance in capturing both chemisorption and physisorption by concurrently optimizing its semilocal component and the long-range van der Waals correction. Internal parameters, including $\kappa$, $b$ in the MS2 functional and $\beta$ in rVV10, were re-fitted concurrently to the adsorption energy of CO/Pt(111) and the Ar2 binding energy, respectively. The resulting Opt(MS+rVV10) functional demonstrates improved accuracy and computational efficiency compared to existing DFAs. Its rigorous parameterization demonstrates a significant improvement over other popular meta-GGA+vdW functionals, such as BEEF-vdW, RPBE+D3, and SCAN+rVV10. This enhancement is evident from its reliable predictions across a broad spectrum of adsorption energies, as illustrated by its performance on the CE39 dataset.
@@
Testing on graphene adsorption on Ni(111), a system known for its challenging double minima in the binding energy curve, revealed that Opt(MS+rVV10) closely aligns with the highly accurate RPA@PBE benchmark and experimental results. Furthermore, additional benchmarking on AE6, BH6, and LC20 sets showed that while Opt(MS+rVV10) offers slightly lower performance for AE6, it remains competitive for hydrogen transfer barriers and lattice constants, delivering overall reliable accuracy across diverse systems. The Opt(MS+rVV10) functional’s improved accuracy in predicting both chemisorption and physisorption energies makes it particularly suited for studying complex catalytic systems such as zeolites, where both strong chemical bonding and weak van der Waals interactions play crucial roles in determining adsorption and reactivity of hydrocarbons. Additionally, its ability to capture subtle electronic interactions suggests potential applications in modeling transition metal catalysts for electrochemical CO2 reduction, where an accurate description of adsorbate binding energies and charge transfer at the electrode-electrolyte interface is critical for predicting catalytic activity and selectivity [72].
IV Computational Details
@@
The Opt(MS+rVV10) functional has been implemented in the developmental version of Vienna Ab-initio Simulation Package (VASP)[73, 74, 75]. All calculations use the pseudopotential project-augmented wave method [76], and a high-energy cutoff of 700 eV to truncate the plane wave basis set. Monkhorst-Pack [77]$k$ meshes for Brillouin zone integration are set to $6\times 6\times 1$ for $2\times 2$ slabs, and $4\times 4\times 1$ for $3\times 3$ slabs and $\Gamma$ point calculation for molecules. Ionic structures were relaxed to a residual force threshold of 0.01 eV/Å per atom with a total energy tolerance of $10^{-5}$eV.
@@
To calculate the binding curves for graphene registry with the Ni(111) surface, we put the graphene sheet on top of the surface (as shown in Fig. 1 of Ref [78]). The metal surface was modeled with a four-layer slab generated with the experimental lattice constants and a 20 Å-thick vacuum layer. The energy cutoff of 600 eV was used, and the $\Gamma$ centered $16\times 16\times 1$ Monkhorst–Pack [77] $k$ meshes were used.
V acknowledgement
This work was supported by the U.S. DOE, Office of Science, Basic Energy Sciences (BES), Grant No. DE-SC0014208. M.K. also acknowledges the support of the American Chemical Society Petroleum Fund, grant number: PRF# 59480-DNI5. Calculations were performed using the National Energy Research Scientific Computing Center (NERSC) using NERSC Awards No. BES-ERCAP0020494 and No. BES-ERCAP0028709 and in part computational resources from Discovery, Research Computing at Northeastern University. A.P. acknowledges support from Delaware Energy Institute, University of Delaware. Q.Z. acknowledges support by Northeastern University, Chemical Engineering Department under start-up funding.
VI Data and Code availability

### v2

@@ 2.3 Determination of parameters for Opt(MS+rVV10)
Here, $\kappa$ and $b$ are tunable parameters in the MS2 functional, while $\beta$ is a parameter in the rVV10 functional. In the original formulation of the MS2 meta-GGA, $\kappa$ = 0.504 and $b$ = 4.0 were optimized to fit the atomization energies of six molecules (AE6 set), and the barrier heights of six reactions (BH6 set) 70. Similarly in rVV10, $\beta$ = 6.3 was fitted to the interaction energies of 22 predominant dispersion-dominated complexes (S22 set) 71.
In this work, we re-optimized $\kappa$, $b$, and $\beta$ to improve the functional’s accuracy for both chemisorption and physisorption. Since the MS2 meta-GGA is known to partially capture intermediate-range vdW interactions 59, it was necessary to re-fit $\beta$ of rVV10 to avoid overestimating vdW forces, which could lead to double-counting. The $\beta$ parameter of rVV10 was therefore optimized for each ( $\kappa$, $b$ ) pair by fitting the binding energy curve of $Ar_{2}$, using highly accurate coupled-cluster CCSD(T) calculations as the reference 72. The $\kappa$ and $b$ of MS2 was then re-optimized by fitting to the experimental adsorption energy of CO on Pt(111), a well-studied surface system with reliable experimental data 73.
We minimized the absolute error in the CO/Pt(111) adsorption energy with respect to $\kappa$ and b using a least-squares fit to a 3rd-order polynomial in the two-dimensional $\kappa$-$b$ space. This approach allowed us to efficiently explore the parameter space and identify the optimal values. The resulting heat map (Figure 2(a)) revealed a well-defined minimum at $\kappa$=0.25 and bb=0.91, with $\beta$ =26.26. This set of parameters delivered the lowest error for the CO/Pt(111) adsorption energy.
When using AE6 atomization energies as the optimization objective for $\kappa$, and $b$, a very different minimum was found, located in the upper-right corner of the $\kappa$-$b$ parameter space (Figure 2(b)). A similar minimum emerged when both the AE6 and BH6 sets were used, in line with the original MS2 parameterization. These findings suggest that developing DFT exchange-correlation functionals for molecular adsorption on solid surfaces requires distinct optimization strategies compared to those used for general chemistry applications. We acknowledge that the CE39 dataset is weighted toward chemisorbed systems, as noted in comprehensive benchmarking studies 39. However, our concurrent optimization approach against both CO/Pt(111) (chemisorption) and Ar2 (physisorption) was specifically designed to address this imbalance and ensure balanced performance across both regimes, as evidenced by our bivariate plot results.
@@
We conducted additional testing of the newly optimized Opt(MS+rVV10) functional on a set of small benchmark systems. These tests included the atomization energies from the AE6 set 70, hydrogen transfer barrier heights from the BH6 set 70, and lattice constants from the LC20 set 74, comprising 20 solids. The performance of Opt(MS+rVV10) was compared against other van der Waals-corrected functionals, including RPBE+D3, BEEF-vdW, and SCAN+rVV10 (the latter being a widely regarded general-purpose functional). The results are summarized in Table 1. For AE6 atomization energies, SCAN+rVV10 and BEEF-vdW showed better accuracy than Opt(MS+rVV10), which exhibited slightly higher errors. For BH6 barrier heights, both BEEF-vdW and Opt(MS+rVV10) performed well, with mean absolute errors (MAEs) of 5.4 kcal/mol and 5.6 kcal/mol, respectively. In the case of LC20 lattice constants, Opt(MS+rVV10) and SCAN+rVV10 were the most accurate, with MAEs of 0.020 Å and 0.021 Å, respectively. These results demonstrate that Opt(MS+rVV10) performs competitively in terms of accuracy, particularly for surface adsorption, while still maintaining reasonable accuracy in other general-purpose tasks.
3 Conclusion
@@
We present the Opt(MS+rVV10), a novel exchange-correlation vdW-corrected functional, representing a significant advancement in the modeling of molecular adsorption on solid surfaces. This functional achieves a balanced performance in capturing both chemisorption and physisorption by concurrently optimizing its semilocal component and the long-range van der Waals correction. Internal parameters, including $\kappa$, $b$ in the MS2 functional and $\beta$ in rVV10, were re-fitted concurrently to the adsorption energy of CO/Pt(111) and the Ar2 binding energy, respectively. The resulting Opt(MS+rVV10) functional demonstrates improved accuracy and computational efficiency compared to existing DFAs. Its rigorous parameterization demonstrates a significant improvement over other popular vdW-corrected functionals, such as BEEF-vdW, RPBE+D3, and SCAN+rVV10. This enhancement is evident from its reliable predictions across a broad spectrum of adsorption energies, as illustrated by its performance on the CE39 dataset.
@@
Testing on graphene adsorption on Ni(111), a system known for its challenging double minima in the binding energy curve, revealed that Opt(MS+rVV10) closely aligns with the highly accurate RPA@PBE benchmark and experimental results. Furthermore, additional benchmarking on AE6, BH6, and LC20 sets showed that while Opt(MS+rVV10) offers slightly lower performance for AE6, it remains competitive for hydrogen transfer barriers and lattice constants, delivering overall reliable accuracy across diverse systems. The Opt(MS+rVV10) functional’s improved accuracy in predicting both chemisorption and physisorption energies makes it particularly suited for studying complex catalytic systems such as zeolites, where both strong chemical bonding and weak van der Waals interactions play crucial roles in determining adsorption and reactivity of hydrocarbons. Additionally, its ability to capture subtle electronic interactions suggests potential applications in modeling transition metal catalysts for electrochemical CO2 reduction, where an accurate description of adsorbate binding energies and charge transfer at the electrode-electrolyte interface is critical for predicting catalytic activity and selectivity 75.
4 Computational Details
@@
The Opt(MS+rVV10) functional has been implemented in the developmental version of Vienna Ab-initio Simulation Package (VASP)76, 77, 78. All calculations use the pseudopotential project-augmented wave method 79, and a high-energy cutoff of 700 eV to truncate the plane wave basis set. Monkhorst-Pack 80kk meshes for Brillouin zone integration are set to $6\times 6\times 1$ for $2\times 2$ slabs, and $4\times 4\times 1$ for $3\times 3$ slabs and $\Gamma$ point calculation for molecules. Ionic structures were relaxed to a residual force threshold of 0.01 eV/Å per atom with a total energy tolerance of $10^{-5}$eV.
@@
To calculate the binding curves for graphene registry with the Ni(111) surface, we put the graphene sheet on top of the surface (as shown in Fig. 1 of Ref 81). The metal surface was modeled with a four-layer slab generated with the experimental lattice constants and a 20 Å-thick vacuum layer. The energy cutoff of 600 eV was used, and the $\Gamma$ centered $16\times 16\times 1$ Monkhorst–Pack 80 $k$ meshes were used.
5 Acknowledgement
This work was supported by the U.S. DOE, Office of Science, Basic Energy Sciences (BES), Grant No. DE-SC0014208. M.K. also acknowledges the support of the American Chemical Society Petroleum Research Fund, grant number: PRF# 59480-DNI5. Calculations were performed using the National Energy Research Scientific Computing Center (NERSC) using NERSC Awards No. BES-ERCAP0020494 and No. BES-ERCAP0028709 and in part computational resources from Discovery, Research Computing at Northeastern University. A.P. acknowledges support from Delaware Energy Institute, University of Delaware. Q.Z. acknowledges support by Northeastern University, Chemical Engineering Department under start-up funding.
6 Data and Code availability